# FAST-GSC: Fast and Adaptive Semantic Transmission for Generative Semantic Communication

### v1

@@ V Performance Evaluation
In this section, we implement the proposed PGSC framework, as well as the two refinement approaches. We conduct extensive experiments that target to answer the following three questions: 1) whether the designed temporal prompt engineering approach can capture the intrinsic sequential dependencies among various semantic units, thereby optimizing the extraction and transmission order to enhance inference performance; and 2) whether the proposed sequential conditional denoising approach can improve the guidance effectiveness of late-arriving semantic units for final performance, thereby mitigating the strict reception immediacy constraints; and 3) whether the two approaches can work synergistically to reduce overall latency while ensuring high task performance. The details of implementations and the analysis of the experimental results are also described.
@@
Experimental Settings: We employ Stable Diffusion v1-$4$https://huggingface.co/CompVis/stable-diffusion-v1-4 as the generative inference model at the receiver and base our optimization strategies on it. As analyzed in Section III-C, the wireless transmission delay is negligible for text prompt transmission. Therefore, we exclude this delay from our experiments. The experiments are conducted on a server with an NVIDIA RTX 4090 GPU with 24 GB of memory. The operating system is Ubuntu 20.04 with Pytorch 2.0.1.
Hyperparameters: We set the total number of denoising steps $M$ at 60. By default, we segment the entire process into intervals of 10 denoising steps and equate the extraction latency of distilling one text unit from the source image to 5 denoising steps. The learning rate is established at 0.009. The entropy coefficient $\xi$ is set at 0.01, the discount factor $\gamma$ is set at 0.99 and the $\eta$ parameter in the clip function is fixed at 0.2.
@@
V-A Effectiveness of Temporal Prompt Engineering
In this part, we evaluate the effectiveness of the proposed temporal prompt engineering (TPE) method. Initially, we demonstrate the training process of our invalid action masking-assisted PPO model and compare its performance with that of two traditional methods. Subsequently, we conduct the statistical analysis of the trained model’s results and investigate the presence of semantic patterns in its decision-making process.
@@ V-A Effectiveness of Temporal Prompt Engineering
Figure 6: Statistical results of the TPE method. (a) Semantic characteristics of text units transmitted in chronological order; (b) Composition of discarded text units; (c) Number of text units transmitted in chronological order. According to our statistical results, six represents the maximum transmission round, as the number of text units transmitted from the seventh transmission onward is zero.
@@
2) Statistic Analysis: Subsequently, we analyze the statistical results of the TPE method. First, we explore the semantic characteristics of text units sent in chronological order, as illustrated in Fig. 6 (a) [^5] Note that to eliminate the impact of different proportions of text unit types in the dataset on the results, we compute the statistics based on their respective proportions over the time series. Consequently, the sum of their proportions at any given moment may exceed 100%.. These results reveal that the TPE method preferentially initiates the transmission of nouns. Consequently, extraction models that target noun retrieval, such as those used for object and scene recognition, should be prioritized in execution. During the second round of transmission, action-related units are more likely to be selected, corresponding to the outputs from action detection models. During the third round of transmission, units related to style and adjectives, which share similar characteristics, are predominantly selected. This indicates that the execution of perceptual extraction models, such as those used for sentiment analysis and style detection, can be appropriately delayed. Units of other types are typically chosen for transmission at later rounds due to their lesser contribution to the semantic information of the inferred images. Additionally, the results indicate that the obtained RL-based TPE approach terminates the extraction process early and discards the transmission of a certain number of text units with a probability of 7.3%. We present the components of the discarded text units in Fig. 6 (b), which shows that the majority of discarded units fall into the ‘OTHERS’ category, such as ‘of’, ‘a’, and ‘the’. This indicates that in the current inference task, these units represent semantic information that is less relevant to the task. Finally, we analyze the number of text units transmitted at different time intervals using the TPE method, as presented in Fig. 6 (c). The results indicate that a large amount of text units are transmitted during the earlier transmission. This is primarily because the semantic guidance effect is most effective at the initial stage of denoising. Consequently, the system prioritizes maintaining a high-performance score, even at the expense of some latency. In the subsequent transmission rounds, the number of transmitted text units decreases to approximately two. During these phases, the extraction latency and segmented denoising duration are closely aligned, indicating the effective parallelization of these processes. In the final transmission rounds, the number of transmitted text units significantly decreases. This reduction is primarily attributed to the smaller number of long prompts in the dataset and the discarding phenomenon of the TPE method.
Figure 7: Visual demonstration and comparison of denoising processes: (a) Conventional GSC, conditioned on the entire prompt throughout the inference process; (b) PGSC, conditioned on text units received in a random sequence over time; (c) TPE-PGSC, conditioned on text units received in a sequence guided by TPE; (d) TPE-PGSC, with changes in an action-related text unit; (e) TPE-PGSC, incorporating a specific style. The residual task latency is defined as the portion of task latency that extends beyond the fixed inference duration and measured in denoising steps.
3) Visual Demonstration and Comparison of Denoising Processes: In Fig. 7, we present a visual demonstration of the denoising processes of the conventional GSC, PGSC, and the TPE-assisted PGSC (TPE-PGSC) approaches. In the conventional GSC framework, the semantic extraction and inference processes are separated. Therefore, the diffusion model can condition on the complete prompt throughout the inference process, as shown in Fig. 7 (a). Although this traditional architecture ensures the quality of inferred images, it also introduces significant latency. In the PGSC framework, although the parallel execution of the semantic extraction and inference processes significantly reduces delays, the quality of the inferred images cannot be guaranteed. As depicted in Fig. 7 (b), the arbitrary assignment of semantic extraction sequences may substantially compromise the quality of inference. In Fig. 7 (c)-(e), we present the denoising processes in the TPE-PGSC framework. By applying lessons learned from RL, we carefully control the sequence in which different categories of semantic units are transmitted, thereby achieving varied degrees of control over the denoising process at different noise levels. This approach ensures that the final inferred image more closely aligns with the original prompt.
V-B Effectiveness of Sequential Conditional Denoising
@@
In this section, we evaluate the performance of the sequential conditional denoising (SCD) method. Initially, we present visual demonstrations of the inputs and outputs from the proposed semantic difference calculation module, showcasing how this module precisely amplifies the modifications required for newly received text units. Subsequently, we provide a visual analysis of the impact of the intervention factor $\alpha$ on the method’s performance.
Figure 8: Detailed visual demonstration of the semantic difference calculation module’s working process. To elucidate the information captured by each feature map, we unfold the conditional noise predictions along the second dimension and visualize the four $64*64$ matrices individually.
@@
1) Visual Demonstration of the Working Process of Semantic Difference Calculation Module: A visual representation of the working processes of the semantic difference calculation module is depicted in Fig. 8. We segment each prompt into two sets, displayed in the first and second columns, respectively. Initially, we transmit the units from the first column to the receiver to initiate the image inference process. Subsequently, after a 20-step denoising period, we transmit the units from the second column. As shown in Fig. 8, direct observation of the predicted noise conditioned on both the previous and the new units may not provide useful information; however, the differences between them reveal significant semantic meanings. This distinction presents the necessary adjustments in the noise space to effectively incorporate the influence of late-arrived units on the inference results. Specifically, in the first transmission scenario, depicted in the first row, the noise difference highlights the edge of the table and the dog’s foot. In the second transmission scenario, the edge of the car, particularly the headlights and tires, is emphasized. This emphasis is attributed to the distinctive, rounded contours characteristic of objects depicted in the Disney style. In the third transmission scenario, the bright areas in the noise difference image represent the woman’s mouth, which is the focal point for achieving the ‘smiling’ directive. By employing this module, we can precisely identify the portions of the noise space that the newly arrived semantic features aim to modify.
@@
2) Visual Demonstration of Various Intervention Values on the Inferred Images: In Fig. 9, we examine the influence of the intervention factor $\alpha$ specified in Equation (14) on the images inferred under different scenarios involving the transmission of various types of semantic units. Similarly, the units from the first column are transmitted to the receiver to commence the image generation process. Following a 20-step denoising period, the units from the second column are subsequently transmitted. We compare the sequential conditional denoising results with intervention factors ranging from $2$ to $20$ against some representative inference failures of the PGSC, which are depicted in the column marked $\alpha=0$. Specifically, in the transmission scenario shown in the first row, when $\alpha=0$, the inferred image resembles an outdoor scene, with wild grass on either side of the dog and the object beneath the dog appearing more like wooden steps. With the increase of $\alpha$, the extended wooden steps transform into a round, enclosed table, and a chair commonly associated with the table appears in the background. Even the wild grass on either side of the dog gradually turns into patterns on the table. In the transmission scenario described in the second row, the image corresponding to $\alpha=0$ depicts the woman with an expression that is more serious than happy. With the increase of $\alpha$, we observe that the woman’s mouth opens wider due to smiling, and her eyes gradually become rounder, displaying an increasingly happy trend. Significantly, starting from $\alpha=10$, we begin to observe the emergence of laugh lines and nasolabial folds on the woman’s face, attributable to the contraction of facial muscles during smiling. In the third row, we investigate the image inference scenario involved with style indications. Although style-related prompts demonstrate greater robustness to the position of the inserted timestamp relative to the total timestamp, from the perspective of color and line contrast, an enhancement in style with the increase of $\alpha$ is still observable. In this last row, we demonstrate the enhancements our proposed method brings to action-related units received later. When $\alpha=0$, the cat’s eyes in the image are wide open, exhibiting no signs of sleepiness. However, as $\alpha$ increases from $2$ to $6$, the cat’s eyes progressively close. At $\alpha=8$, the cat’s open eyes and its facial patterns begin to blend together. From $\alpha=10$ onward, the cat’s eyes are clearly closed, with the area previously occupied by the eyes now resembling the pattern of eyelids. Notably, at $\alpha=18$ and $\alpha=20$, the cat’s ears are observed to droop, indicative of drowsiness. However, it is important to note that these visual results also demonstrate that a higher $\alpha$ is not always better. We observe that when $\alpha$ is very high, the semantic information related to previously received units is diminished in the inferred images. For example, at $\alpha=20$, a strange pattern appears on the dog’s nose in the first column, and the aesthetics of the woman’s image in the second column are also reduced. Therefore, in practice, it is necessary to design an appropriate intervention value to balance the influence of previously received units and newly received units on the final inferred image.
V-C Overall Analysis of Combined Refinements
In this section, we investigate whether the TPE and SCD methods can work together to enhance the system performance of the PGSC framework. Initially, we explore the impact of various intervention factors of the SCD method to find a more suitable one for use with the TPE method under different extraction latency settings. Following this, we synergistically integrate these enhancements into the PGSC framework, ultimately establishing the FAST-GSC framework. Subsequently, we present numerical results, comparing the conventional GSC framework with the PGSC framework featuring different enhancements, and the FAST-GSC framework.
@@ V-C Overall Analysis of Combined Refinements
1) Numerical Results of Various Intervention Factor Values on the System Performance: Herein, we examine the variations in CLIP scores with respect to the intervention factor $\alpha$, aiming to identify a more suitable value to work with TPE methods under our default and two alternative extraction latency settings. To synchronize with the training process of the TPE method, we segment every 5 denoising steps for the alternative setting where the latency for extracting a single semantic unit ${\tau_{e}}$ equals the duration of 2.5 denoising steps, every 10 denoising steps for the default setting where the latency for extracting a single semantic unit ${\tau_{e}}$ equals the duration of 5 denoising steps, and every 15 denoising steps for the alternative setting where the latency for extracting a single semantic unit ${\tau_{e}}$ equals the duration of 7.5 denoising steps. As illustrated in Fig. 10, under the same denoising speed, a higher intervention value is desired for systems with longer extraction latency. This is due to the fact that, once the initial semantic units initiate the denoising process, the extended extraction latency prolongs the incorporation time for each subsequent semantic unit. Consequently, the residual noise level in intermediate results is lower compared to settings with shorter extraction latency. In such scenarios, a greater noise space intervention capability is necessary to adjust the noise space effectively and incorporate the newly arrived units under conditions of extended extraction latency.
Figure 11: Numerical results: (a) Task performance, measured by CLIP score; (b) Residual task latency, defined as the portion of task latency that extends beyond the fixed inference duration and measured in denoising steps; (c) Task efficiency, defined as the ratio of the task performance to the latency count.
2) Numerical Results of Task Performance and Latency: Herein, we focus on our default extraction speed setting, where the extraction latency for distilling one text unit from the source image corresponds to 5 denoising steps. Based on the numerical results assessing the impact of various intervention factor values on system performance, we set the intervention factor $\alpha$ at 4 to synergistically integrate the TPE and SCD methods to establish the FAST-GSC framework. Finally, we present the task performance, latency count and task efficiency results of the conventional GSC, PGSC with random transmission order, TPE-PGSC, SCD-PGSC, and FAST-GSC in Fig. 11. Herein, task efficiency is defined as the ratio of task performance to residual task latency. Numerical results indicate that, although the direct implementation of PGSC effectively reduces system latency, it leads to significant performance degradation. Enhanced by the TPE method, the TPE-PGSC framework further reduces the residual task latency of PGSC by 21% and boosts the CLIP score by 26%. Additionally, the SCD-PGSC framework improves the CLIP score by 8% over the direct PGSC implementation. In terms of task efficiency, the FAST-GSC framework emerges as the most beneficial. It achieves a performance score comparable to the conventional GSC architecture while realizing a 52% reduction in residual task latency.

### v2

@@ V Performance Evaluation
In this section, we implement the proposed FAST framework. The details of implementations and the analysis of the experimental results are also described.
@@
Experimental Settings: We employ Stable Diffusion v1-4444https://huggingface.co/CompVis/stable-diffusion-v1-4 as the generative reconstruction model at the receiver and base our optimization strategies on it. The experiments are conducted on a server with an NVIDIA RTX 4090 GPU with 24 GB of memory. The operating system is Ubuntu 20.04 with Pytorch 2.0.1. We considered an extreme scenario where the computational resources at the edge are severely constrained, allowing for the execution of merely a single semantic extraction model per user at any given time. In the multi-user transmission setting, 5 users are uniformly distributed within a 10-meter radius of an edge server. The task importance scores assigned to the users are 5, 3, 1, 1, and 1, respectively. The large-scale fading is modeled as (35.3+37.6log10(d))-(35.3+37.6*log10(d)) in dB [56], where $d$ is the distance between a user and the edge server in meters. For small-scale fading, Rayleigh fading model is adopted. The transmitting power and noise power are specified as $10$ dBm and 80-80 dBm, respectively. The maximum tolerable communication delay is set to 0.01 seconds. We assume the packet size is 256 bits for every semantic unit. The actual transmission rate is determined by the selected modulation and coding scheme [57].
Training Hyperparameters: We set the total number of denoising steps $M$ at 60. By default, we segment the entire process into intervals of 10 denoising steps and equate the extraction latency of distilling one text unit from the source image to 5 denoising steps. For the training of temporal prompt engineering and semantic-aware resource allocation modules, the learning rates are set to 0.009 and 0.00015, with entropy coefficients $\xi$ of 0.01 and 0.03, respectively. A shared discount factor $\gamma$ of 0.99 and a clip parameter $\eta$ of 0.2 are used for both.
@@
V-A Evaluation of the Temporal Prompt Engineering Module
In this part, we evaluate the effectiveness of the proposed temporal prompt engineering module. Initially, we demonstrate the training process of our invalid action masking-assisted PPO model and compare its performance with that of two traditional methods. Subsequently, we conduct the statistical analysis of the trained model’s results and investigate the presence of semantic patterns in its decision-making process.
@@ V-A Evaluation of the Temporal Prompt Engineering Module
Figure 7: Statistical results of the temporal prompt engineering method. (a) Semantic characteristics of text units transmitted in chronological order; (b) Composition of discarded text units; (c) Number of text units transmitted in chronological order. According to our statistical results, six represents the maximum transmission round, as the number of text units transmitted from the seventh transmission onward is zero.
@@
2) Statistic Analysis: Subsequently, we analyze the statistical results of the temporal prompt engineering module. First, we explore the semantic characteristics of text units sent in chronological order, as illustrated in Fig. 7 (a) [^5] Note that to eliminate the impact of different proportions of text unit types in the dataset on the results, we compute the statistics based on their respective proportions over the time series. Consequently, the sum of their proportions at any given moment may exceed 100%.. These results reveal that the temporal prompt engineering module preferentially initiates the transmission of nouns. Consequently, extraction models that target noun retrieval, such as those used for object and scene recognition, should be prioritized in execution. During the second round of transmission, action-related units are more likely to be selected, corresponding to the outputs from action detection models. During the third round of transmission, units related to style and adjectives, which share similar characteristics, are predominantly selected. This indicates that the execution of perceptual extraction models, such as those used for sentiment analysis and style detection, can be appropriately delayed. Units of other types are typically chosen for transmission at later rounds due to their lesser contribution to the semantic information of the inferred images. Additionally, the results indicate that the obtained RL-based temporal prompt engineering module terminates the extraction process early and discards the transmission of a certain number of text units with a probability of 7.3%. We present the components of the discarded text units in Fig. 7 (b), which shows that the majority of discarded units fall into the ‘OTHERS’ category, such as ‘of’, ‘a’, and ‘the’. This indicates that in the current reconstruction task, these units represent semantic information that is less relevant to the task. Finally, we analyze the number of text units transmitted at different time intervals adopting the temporal prompt engineering module, as presented in Fig. 7 (c). The results indicate that a large amount of text units are transmitted during the earlier transmission. This is primarily because the semantic guidance effect is most effective at the initial stage of denoising. Consequently, the system prioritizes maintaining a high-performance score, even at the expense of some latency. In the subsequent transmission rounds, the number of transmitted text units decreases to approximately two. During these phases, the extraction latency and segmented denoising duration are closely aligned, indicating the effective parallelization of these processes. In the final transmission rounds, the number of transmitted text units significantly decreases. This reduction is primarily attributed to the smaller number of long prompts in the dataset and the discarding phenomenon of the temporal prompt engineering module.
V-B Evaluation of the Sequential Conditional Denoising Module
@@
In this section, we evaluate the performance of the sequential conditional denoising module. Initially, we present visual demonstrations of the inputs and outputs from the proposed semantic difference calculation subunit, showcasing how this subunit precisely amplifies the modifications required for newly received text units. Subsequently, we provide a visual analysis of the impact of the intervention factor $\alpha$ on the method’s performance.
Figure 8: Detailed visual demonstration of the semantic difference calculation module’s working process. For each row, the intermediate output of a representative image generation is presented. The visualization is achieved by unfolding the latent representation of the conditional noise predictions across the channel dimension (4 channels), revealing the distinct information embedded in each 646464*64 feature map.
@@
1) Visual Demonstration of the Working Process of Semantic Difference Calculation Module: A visual representation of the working processes of the semantic difference calculation module is depicted in Fig. 8. We segment each prompt into two sets, displayed in the first and second columns, respectively. Initially, we transmit the units from the first column to the receiver to initiate the image reconstruction process. Subsequently, after a 20-step denoising period, we transmit the units from the second column. As shown in Fig. 8, direct observation of the predicted noise conditioned on both the previous and the new units may not provide useful information; however, the differences between them reveal significant semantic meanings. This distinction presents the necessary adjustments in the noise space to effectively incorporate the influence of late-arrived units on the reconstruction results. Specifically, in the first transmission scenario, depicted in the first row, the noise difference highlights the edge of the table and the dog’s foot. In the second transmission scenario, the edge of the car, particularly the headlights and tires, is emphasized. This emphasis is attributed to the distinctive, rounded contours characteristic of objects depicted in the Disney style. In the third transmission scenario, the bright areas in the noise difference image represent the woman’s mouth, which is the focal point for achieving the ‘smiling’ directive. By employing this module, we can precisely identify the portions of the noise space that the newly arrived semantic features aim to modify.
@@
2) Visual Demonstration of Various Intervention Values on the Inferred Images: In Fig. 9, we examine the influence of the intervention factor $\alpha$ specified in Equation (11) on the images inferred under different scenarios involving the transmission of various types of semantic units. Similarly, the units from the first column are transmitted to the receiver to commence the image generation process. Following a 20-step denoising period, the units from the second column are subsequently transmitted. We compare the reconstruction results using sequential conditional denoising with intervention factors ranging from $2$ to $20$ against representative failure cases produced by the traditional denoising method (shown in the column labeled $\alpha=0$). Take the transmission scenario in the first row as an example. When $\alpha=0$, the inferred image resembles an outdoor scene, with wild grass on either side of the dog and the object beneath the dog appearing more like wooden steps. With the increase of $\alpha$, the extended wooden steps transform into a round, enclosed table, and a chair commonly associated with the table appears in the background. Even the wild grass on either side of the dog gradually turns into patterns on the table. However, it is important to note that these visual results also demonstrate that a higher $\alpha$ is not always better. We observe that when $\alpha$ is very high, the semantic information related to previously received units is diminished in the inferred images. For example, at $\alpha=20$, a strange pattern appears on the dog’s nose in the first column, and the aesthetics of the woman’s image in the second column are also reduced. Therefore, in practice, it is necessary to design an appropriate intervention value to balance the influence of previously received units and newly received units on the final inferred image.
V-C Evaluation of the Semantic-aware Resource Allocation Module
Figure 10: Validation reward relative to the training iterations. In (a), the available bandwidth is 10 KHz. In (b), the available bandwidth is 20 KHz.
In this section, we evaluate the performance of the semantic-aware resource allocation module. The proposed method is compared against the following two benchmark policies: 1) Random Policy: The available bandwidth is randomly allocated among all extracted semantic units. 2) Baseline Policy: The limited bandwidth is allocated based on user task importance, prioritizing users with higher task importance. The results are shown in Fig. 10 (a) and (b), where the available bandwidth is $10$ KHz and $20$ KHz, respectively. In Fig. 10 (a), the proposed method’s mean reward rises until converging around episode 580, stabilizing between 15–16. In contrast, the Baseline Policy remains below 12, and the Random Policy performs worst, staying below 6. In Fig. 10 (b), as the bandwidth constraint is relaxed, the mean reward attained by the proposed method continues to increase until convergence occurs at approximately 1000 episodes, stabilizing at a value between 18 and 19. Meanwhile, the mean reward of the Baseline Policy is around 15, whereas the Random Policy remains consistently low at around 13.
Across both Fig. 10 (a) and (b), the proposed method consistently outperforms both the Baseline Policy and the Random Policy in terms of absolute reward value. This result demonstrates its strong adaptability to channel state variations and its overall superiority in optimizing resource allocation.
V-D Overall Analysis of Combined Refinements
In this section, we examine whether the proposed three modules can function cohesively to improve overall system performance. Initially, we explore the impact of various intervention factors of the sequential conditional denoising module to find a more suitable one for use in the FAST framework. Subsequently, we present numerical results comparing the system precision, latency, and efficiency of the proposed FAST framework against those of conventional GSC systems.
@@ V-D Overall Analysis of Combined Refinements
1) Numerical Results of Various Intervention Factor Values on the System Performance: Herein, we examine the variations in CLIP scores with respect to the intervention factor $\alpha$, aiming to identify a more suitable value to work within the FAST under our default and two alternative extraction latency settings. To synchronize with the training process of the temporal prompt engineering method, we segment every 5 denoising steps for the alternative setting where the latency for extracting a single semantic unit $\tau_{e}$} equals the duration of 2.5 denoising steps, every 10 denoising steps for the default setting where the latency for extracting a single semantic unit $\tau_{e}$} equals the duration of 5 denoising steps, and every 15 denoising steps for the alternative setting where the latency for extracting a single semantic unit $\tau_{e}$} equals the duration of 7.5 denoising steps. As illustrated in Fig. 11, under the same denoising speed, a higher intervention value is desired for systems with longer extraction latency. This is due to the fact that, once the initial semantic units initiate the denoising process, the extended extraction latency prolongs the incorporation time for each subsequent semantic unit. Consequently, the residual noise level in intermediate results is lower compared to settings with shorter extraction latency. In such scenarios, a greater noise space intervention capability is necessary to adjust the noise space effectively and incorporate the newly arrived units under conditions of extended extraction latency.
Figure 12: Comparison of numerical results for the conventional GSC scheme, the full FAST framework, and its variants incorporating temporal prompt engineering (TPE), semantic-aware resource allocation (SRA), and sequential conditional denoising (SCD) modules. (a) System precision: the weighted average of individual reconstruction precision; (b) Residual system latency: the weighted average of individual residual latency, which is defined as the latency exceeding the fixed reconstruction duration and measured in denoising steps; (c) System efficiency: the weighted average of individual efficiency, which is defined as the ratio of a user’s precision score to their residual latency.
2) Numerical Results of System Performance: Based on the numerical results assessing the impact of various intervention factor values on the reconstruction precision, we set the intervention factor $\alpha$ at 4. To comprehensively evaluate the proposed framework, we define three performance metrics for each user: precision, residual latency, and efficiency. Precision quantifies reconstruction fidelity and is measured by the CLIP score. Residual latency is the latency exceeding the fixed reconstruction duration, measured in denoising steps. Efficiency is defined as the ratio of a user’s precision to their residual latency. The corresponding system-level metrics are then derived as the weighted average of individual user metrics, weighted by each user’s task importance score. Finally, a comprehensive evaluation of the conventional GSC scheme, FAST variants, and the full FAST framework is presented in Fig. 12. Numerical results demonstrate that integrating the temporal prompt engineering and semantic-aware resource allocation modules into the FAST framework collectively reduces residual system latency by 21% and improves system precision by 29%, compared to a baseline without these components. Furthermore, the sequential conditional denoising module contributes an additional 8% gain in system precision. In terms of overall efficiency, the FAST framework achieves a metric approximately 1.6 times higher than that of conventional GSC. These findings confirm that the proposed architecture exhibits enhanced potential for deployment in multi-user GSC scenarios characterized by stringent communication and computational constraints.